OpenStudy (thats_kyy):

A husband and wife attend the local recreation center. The husband pays a flat rate of $125 for one year, shown as f(x). The wife pays $6 per visit, shown as g(x). Which function shows the correct combination of these two functions to represent the total cost to them of attending the recreation center in one year, shown as h(x)? f(x) = 125x, g(x) = 6, h(x) = 125x + 6 f(x) = 125, g(x) = 6, h(x) = 6 + 125 f(x) = 125x, g(x) = 6x, h(x) = 125x + 6x f(x) = 125, g(x) = 6x, h(x) = 6x + 125

OpenStudy (thats_kyy):

@triciaal

OpenStudy (thats_kyy):

@Nnesha

OpenStudy (thats_kyy):

@Loser66 @texaschic101

Nnesha (nnesha):

cost is represent by y variable and there is a key word `per` the number with per word represent slope and the slope would be with x variable (y=mx+b) lol

OpenStudy (thats_kyy):

Oh

OpenStudy (thats_kyy):

okay now i feel stupid -_-

Nnesha (nnesha):

hmm why didn't understand that ? it's okay

OpenStudy (thats_kyy):

idk because all i see is numbers and letters and its confusing -_-

Nnesha (nnesha):

well f(x) = cost which is what ? `The husband pays a flat rate of $125 for one year, shown as f(x). `

OpenStudy (thats_kyy):

so f= the flat rate

Nnesha (nnesha):

rate is what ?? in numbers

OpenStudy (thats_kyy):

125$

OpenStudy (thats_kyy):

@texaschic101

OpenStudy (thats_kyy):

@Nnesha

Nnesha (nnesha):

right so f(x) =125 which is same as y=125 like i said y represent cost

Nnesha (nnesha):

now next one

Nnesha (nnesha):

`The wife pays $6 per visit, shown as g(x).` remember what i told you about `per` word write the equation g(x) = ?

OpenStudy (thats_kyy):

g(x) = 6 and per means multiply

Nnesha (nnesha):

yes right so use the x variable

OpenStudy (thats_kyy):

x= 2

Nnesha (nnesha):

we don't know how many visits

Nnesha (nnesha):

so we don't know what x is just write the variable

OpenStudy (thats_kyy):

ok

Nnesha (nnesha):

so g(x) = ??

OpenStudy (thats_kyy):

6?

Nnesha (nnesha):

u said per means multiply so $6 per visits (we don't know how many visit so use x variable for number of visit ) then how would you write `$6 per x`

OpenStudy (thats_kyy):

6(x)

Nnesha (nnesha):

right same as 6x

OpenStudy (thats_kyy):

yep

Nnesha (nnesha):

now add both functions to find h(x) \[\huge\rm h(x)=f(x)+g(x) \]

OpenStudy (thats_kyy):

125x

OpenStudy (thats_kyy):

i think i got it B?

Nnesha (nnesha):

what we got for f(x) and for g(x) f(x) = ?? g(x) = ?

OpenStudy (thats_kyy):

6xand 125

Nnesha (nnesha):

right so replace f(X) with 125 and g(x) with 6x \[\huge\rm h(x)=f(x)+g(x) \]

OpenStudy (thats_kyy):

ok hold on

OpenStudy (thats_kyy):

OOOOO I got it I got Nesha!!

Nnesha (nnesha):

*nnesha so which one is correct ?? :=)

OpenStudy (thats_kyy):

im stuck in between D and C but i think D

Nnesha (nnesha):

why do you think its C ?? you know f(x) = 125 and g(x) = 6x

OpenStudy (thats_kyy):

Because of the x's but i guess not

Nnesha (nnesha):